OpenStudy (anonymous):

if I have the hypotenuse of a right triangle, do I have to square it when trying to find the legs using the pythagorean theorem

OpenStudy (jdoe0001):

yes

OpenStudy (solomonzelman):

\[a^2+b^2=c^2\] solving for hypotenuse (or for c), then use, \[\sqrt{ a^2+b^2}=c\]

OpenStudy (anonymous):

thank you